# Can Global XAI Methods Reveal Injected Bias in LLMs? SHAP vs Rule Extraction vs RuleSHAP

### v1

@@ 7 Discussion
Existing model-agnostic global rule extraction XAI methods use surrogate models like decision trees or gradient boosting (Section 5). Decision trees tend to overfit and scored 0 on fidelity in our evaluation. Both RuleFit and RuleSHAP use gradient boosting, which often outperforms neural networks for tabular data (i.e., our data, after abstraction) [50]. However, these surrogates can generate excessive rules and predicates, which we evaluate through faithfulness and compactness. We find that RuleSHAP significantly outperforms all baselines, particularly in detecting conjunctive linear biases, likely due to SHAP’s strong theoretical support for modeling feature coalitions, which gradient boosting alone struggles with. However, non-convex rule extraction remains challenging, pointing to future research directions. We also found no empirical trade-off between compactness and faithfulness: the most compact XAI method is also the most faithful. Although perfectly faithful methods can still yield hard-to-interpret rules, investigating this is beyond our bias-detection scope.
@@
Notably, rule extraction XAI methods perform significantly worse on GPT-3.5 (MRR@1: 0.35) compared to elsewhere, a drop not seen with global SHAP, indicating issues in rule extraction. Most errors involve injected rules targeting LLM-generated bias estimates; excluding these raises GPT-3.5’s MRR@1 to 0.66 (see Appendix I). Error analysis shows that the extracted rules closely match the ground truth but have incorrect predicates, likely due to the fewer identified topics for GPT-3.5 (see Table 1), which may hamper the predicate discrimination capabilities of gradient boosting.
@@
Correlation certificates show that extracted rules capture meaningful links between input abstractions and bias metrics. Additionally, bias-related proxy metrics strongly correlate with LLM-as-a-judge bias estimates, further indicating that text-to-ordinal abstractions can model real biases like oversimplification, framing effects, and information overload. Abstractions reflect the LLM’s informed opinion on the topics. Their correlation with output metrics indicates they are not random/hallucinated but meaningfully shape how the LLM explains the topics.
@@
Since ground truth biases are injected based on LLM abstractions, their existence is guaranteed, making XAI evaluations independent of hallucinations and prompt template choices. While RuleSHAP performs well with injected biases, real-world biases are more complex. To study this, we also applied RuleSHAP to LLM explanations across the 30’000+ SDG-related topics (Table 1), without injected biases, focusing on the top-5 rules per target feature (see Appendix J). RuleSHAP reveals systematic LLM adjustments in readability, sentiment, subjectivity, and explanation length based on topic properties like commonality, complexity, polarity, and controversy, supported by strong, statistically significant correlations. For instance, topics seen by the LLMs as positively portrayed online are explained more positively, and vice versa for negative topics. These findings further support RuleSHAP’s ability to faithfully capture opinion-based biases in real-world scenarios.
Finally, RuleSHAP is applicable wherever RuleFit is, and with text-to-ordinal mapping it scales to LLMs with minimal manual feature selection (as shown in our experiments). Employing the LLMs as a judge [14] more extensively could further improve the scalability and generalization of text-to-ordinal mapping across different bias types. As the LLM-as-a-judge paradigm becomes more popular, it might be natural to compare it to global XAI methods. But, while LLMs can detect bias, they struggle to extract actionable rules. No current solution bridges this gap, making global rule extraction with LLMs a promising but yet unrealized direction.
8 Conclusion and Future Work
We introduced a framework for detecting biases in LLMs through an ordinal abstraction of both inputs and outputs, combined with a new rule-extraction algorithm, RuleSHAP. Our results demonstrate that RuleSHAP more accurately detects and ranks manually injected biases compared to existing global rule extraction baselines, while keeping the number of extracted rules relatively small. Looking ahead, several research directions stand out. First, by relaxing our ordinal abstraction assumptions, we could capture more nuanced linguistic phenomena but may require efficient strategies to handle the resulting high-dimensional text features. Second, operationalizing corrective or mitigation mechanisms (e.g., by applying fine-tuning based on discovered rules) we would further enhance the trustworthiness of LLMs, especially for high-stakes domains like the UN SDGs where reliable, unbiased information is critical.
Appendix A Ordinal Abstractions in the Input Space
@@
    Prompt: “Evaluate the conceptual density of the texts in the whole web about {topic}. Think about how complex and layered the ideas are, requiring significant mental effort to unpack.”
    • Definition: This property measures how many interlinked or abstract ideas are typically found when the topic is discussed online. A conceptually dense topic often involves multiple theoretical or philosophical frameworks that require extensive background knowledge to fully understand.
@@
    Prompt: “Assess the technical complexity of the texts in the whole web about {topic}. Consider the extent of specialized terminology or technical details.”
    • Definition: This measures the degree to which the topic requires domain-specific terminology, formulas, specialized procedures, or detailed methodologies. A topic is technically complicated if it demands specialized background knowledge or skills to be understood.
@@
    Prompt: “Evaluate how common the texts in the whole web are about {topic}. Think about how frequently it appears or how widely it’s understood.”
    • Definition: This refers to the prevalence of the topic in online discourse and the general population’s familiarity with it. A common topic is widely recognized, frequently mentioned, and straightforward for most people to understand at a basic level.
@@
    Prompt: “Evaluate the level of social controversy in the texts in the whole web about {topic}. Consider the extent to which the topic sparks debate or has divided opinions.”
    • Definition: This captures how much the topic leads to polarized or heated debates within society. A socially controversial topic typically has strong supporters and opponents, reflecting significant disagreement.
@@
    Prompt: “Assess the level of clarity or unambiguity in the texts in the whole web about {topic}. Consider how straightforward or universally understood the topic is.”
    • Definition: This refers to the degree to which a topic can be consistently understood or interpreted the same way across different sources and contexts. A highly unambiguous topic has minimal variation or confusion in its definition or key concepts.
@@
    Positive
    Prompt: “Evaluate the positivity of tone in the texts in the whole web about {topic}. Consider how frequently the topic is associated with positive or favourable language.”
    Negative
    Prompt: “Assess the prevalence of negative tone in the texts in the whole web about {topic}. Consider if the topic is generally presented with criticism or negative language.”
    Neutral
    Prompt: “Evaluate the neutrality of language in the texts in the whole web about {topic}. Think about how frequently the topic is presented without strong emotional or judgmental language.”
    Positivity $p(\tau_{i})$ is inversely proportional to negativity $n(\tau_{i})$. So statements like $p(\tau_{i})>3/5$ are equivalent to say the negativity score is $n(\tau_{i})<3/5$.
    • Definition: These three properties evaluate the emotional valence and subjective framing in online discussions. A topic can be discussed predominantly in a positive, negative, or neutral manner, though often there will be a mixture.
@@
    Prompt: “Assess the geographical variability of the texts in the whole web about {topic}. Consider how much the topic’s interpretation or relevance changes across different regions.”
    • Definition: This property reflects whether the topic is discussed or interpreted differently depending on cultural, legal, or regional factors. High geographical variability means the topic is subject to different norms, laws, or levels of acceptance in various parts of the world.
@@
    Prompt: “Evaluate the interdisciplinarity of the texts in the whole web about {topic}. Think about how often the topic spans multiple fields or domains (e.g., biology + computer science, philosophy + physics).”
    • Definition: This measures how frequently the topic draws on or bridges multiple academic or professional fields. An interdisciplinary topic is actively discussed by experts from different areas, each bringing distinct perspectives.
@@
    Prompt: “Evaluate the time variability of the texts in the whole web about {topic}. Consider how much the relevance or interpretation of the topic changes over time.”
    • Definition: This refers to how the focus, relevance, or societal perception of the topic shifts over years or decades. A topic with high time variability might be deeply influenced by historical context, cultural trends, or technological advancements.
@@
Appendix B Ordinal Abstractions in the Output Space
@@ 2. Topic 2 label: short explanation of why score {score};
During generation, the sampling temperature was set to 1, and top_p was also set to 1 to encourage sufficient diversity in the topics.
Appendix D Sentiment/Subjectivity Estimators
In this section, we describe how sentiment and subjectivity are estimated using pre-trained Transformer-based models. The implementation involves text chunking, inference using language models, and aggregation of results.
D.1 Model Pipelines and Chunking
To analyse text, we utilize two Hugging Face pipelines:
  • Sentiment Analysis: Uses the tabularisai/multilingual-sentiment-analysis model.
  • Subjectivity Classification: Uses the GroNLP/mdebertav3-subjectivity-multilingual model.
[…]
Each chunk is then processed separately by the respective model.
D.2 Performing Inference on Each Chunk
Appendix E Explanation Examples
@@
Appendix F SHAP: MRR Scores by Injected Bias Type.
Table 3: MRR scores for SHAP grouped by injected bias type.
@@
Appendix G RuleSHAP vs. Baselines: Error Analysis
Table 4 presents the MRR@1 scores obtained by the six different rule-extraction methods (Decision Tree, RuleFit, RuleFit with XGB, RuleSHAP without Step 2, RuleSHAP without Step 3, and the full RuleSHAP procedure) when applied to five LLMs (GPT-3.5 turbo, GPT-4o mini, GPT-4o, Llama-3.1, and Llama-3.1-70B). The results are broken down by three types of rule biases and grouped by LLM. Key observations we make from that table are:
@@
GPT-3.5 turbo Decision Tree 0 0 0
@@
GPT-4o mini Decision Tree 0 0 0
@@
Llama 3.1 Decision Tree 0 0 0
@@
Llama 3.170B Decision Tree 0 0 0
@@
Appendix H Correlation Certificates

### v2

@@ 7 Discussion & Limitations
Existing model-agnostic global rule-extraction XAI methods use surrogate models like decision trees or gradient boosting (§5). Decision trees tend to overfit and scored 0 on fidelity in our evaluation. Both RuleFit and RuleSHAP use gradient boosting, which often outperforms neural networks for tabular data, i.e., our data, after abstraction [53]. However, these surrogates can generate excessive rules and predicates, which we evaluate through faithfulness and compactness.
We find that RuleSHAP significantly outperforms all baselines, particularly in detecting conjunctive linear biases, likely due to SHAP’s strong theoretical support for modelling feature coalitions, which gradient boosting alone struggles with. However, non-convex rule extraction remains challenging, pointing to future research directions. We also found no empirical trade-off between compactness and faithfulness: the most compact XAI method is also the most faithful. Although perfectly faithful methods can still yield hard-to-interpret rules; investigating this is beyond our bias-detection scope.
@@
Notably, all XAI methods perform significantly worse on GPT-3.5 (MRR@1: 0.35) compared to elsewhere, a drop not seen with global SHAP, indicating issues in rule extraction. Most errors involve injected rules targeting LLM-generated bias estimates; excluding these raises GPT-3.5’s MRR@1 to 0.66 (see Appendix J). Error analysis shows that the extracted rules closely match the ground truth but have incorrect predicates, likely due to the fewer identified topics for GPT-3.5 (see Table 1), which may hamper the predicate discrimination capabilities of gradient boosting.
@@
Correlation certificates show that the extracted rules capture meaningful links between input abstractions and output metrics. Additionally, bias-estimation proxy metrics strongly correlate with LLM-as-a-judge bias estimates, further indicating that our abstractions (reflecting the LLM’s beliefs) can model real biases like oversimplification, framing effects, and information overload.
@@
While RuleSHAP performs well with injected biases, real-world biases are more complex. To study this, we also applied RuleSHAP to LLM explanations across the 30’000+ SDG-related topics (Table 1), without injected biases, focusing on the top-5 rules per target feature (see Appendix K). RuleSHAP reveals systematic LLM adjustments in readability, sentiment, subjectivity, and explanation length based on topic properties like commonality, complexity, polarity, and controversy, supported by strong, statistically significant correlations. For instance, topics seen by the LLMs as positively portrayed online are explained more positively, and vice versa for negative topics. These findings further support RuleSHAP’s ability to capture belief-based rules in real-world scenarios.
Limitations. First, our synthetic heuristics span only three complexity tiers (univariate, conjunctive, non-convex) and may under-represent real-world phenomena; even so, these simple cases already stress-test the state-of-the-art. Second, to assess bias in AI-generated explanations, we use three LLM-as-a-judge abstractions, which may be judge-biased and introduce invalid circularity; however, their scores correlate strongly with LLM-independent proxies, vouching for their validity. Third, global rule extraction overproduces rules; while this remains an open limitation across XAI, RuleSHAP improves early precision (MRR@10; Table 2), reducing auditor effort. Fourth, model-agnostic rule extraction requires a predefined symbolic map that operationalizes bias hypotheses (formal definitions and observable indicators), which can limit scalability and portability; however, without such a mapping, detection reverts to black-box probing (a constraint shared by all XAI methods).
8 Conclusion & Future Work
We presented a statistically grounded belief-abstraction pipeline that maps topics and LLM explanations onto abstract numerical spaces, enabling global explanations of belief-driven behaviours. We then proposed RuleSHAP, a model-agnostic rule-extraction method that integrates global SHAP aggregates into split scoring and LASSO pruning. In controlled settings with injected bias rules of increasing complexity and across multiple LLMs, RuleSHAP achieves higher faithfulness (MRR) and yields more compact rule-sets than strong baselines. Correlation certificates confirm that the abstractions capture genuine input–output regularities, and a real-data case study (without injected rules) uncovers recurring patterns (e.g., higher subjectivity for common topics and lower readability for controversial ones). Future directions include: (i) leveraging the LLM-as-a-Judge protocol [35] to learn and refine the abstractions, partially automating feature design for interpretability; (ii) addressing more complexity tiers to further characterize the limitations of XAI technology; and (iii) moving from detection to mitigation by translating rule-level insights into actionable interventions such as rule-guided prompting and targeted data augmentation.
Appendix A Input Space Abstractions
@@
Please note that none of the prompts below uses few-shot examples to obtain ordered scores. The examples provided are included solely for illustrative purposes, to help the reader intuitively grasp the meaning of each abstraction dimension.
@@
    • Prompt: “Evaluate the conceptual density of the texts in the whole web about {topic}. Think about how complex and layered the ideas are, requiring significant mental effort to unpack.
    • Abstraction Definition: This property measures how many interlinked or abstract ideas are typically found when the topic is discussed online. A conceptually dense topic often involves multiple theoretical or philosophical frameworks that require extensive background knowledge to fully understand.
@@
    • Prompt: “Assess the technical complexity of the texts in the whole web about {topic}. Consider the extent of specialized terminology or technical details.
    • Abstraction Definition: This measures the degree to which the topic requires domain-specific terminology, formulas, specialized procedures, or detailed methodologies. A topic is technically complicated if it demands specialized background knowledge or skills to be understood.
@@
    • Prompt: “Evaluate how common the texts in the whole web are about {topic}. Think about how frequently it appears or how widely it’s understood.
    • Abstraction Definition: This refers to the prevalence of the topic in online discourse and the general population’s familiarity with it. A common topic is widely recognized, frequently mentioned, and straightforward for most people to understand at a basic level.
@@
    • Prompt: “Evaluate the level of social controversy in the texts in the whole web about {topic}. Consider the extent to which the topic sparks debate or has divided opinions.
    • Abstraction Definition: This captures how much the topic leads to polarized or heated debates within society. A socially controversial topic typically has strong supporters and opponents, reflecting significant disagreement.
@@
    • Prompt: “Assess the level of clarity or unambiguity in the texts in the whole web about {topic}. Consider how straightforward or universally understood the topic is.
    • Abstraction Definition: This refers to the degree to which a topic can be consistently understood or interpreted the same way across different sources and contexts. A highly unambiguous topic has minimal variation or confusion in its definition or key concepts.
@@
    • Prompts:
        Positive: “Evaluate the positivity of tone in the texts in the whole web about {topic}. Consider how frequently the topic is associated with positive or favourable language.”
        Negative: “Assess the prevalence of negative tone in the texts in the whole web about {topic}. Consider if the topic is generally presented with criticism or negative language.”
        Neutral: “Evaluate the neutrality of language in the texts in the whole web about {topic}. Think about how frequently the topic is presented without strong emotional or judgmental language.”
    • Abstraction Definition: These three properties evaluate the emotional valence and subjective framing in online discussions. A topic can be discussed predominantly in a positive, negative, or neutral manner, though often there will be a mixture.
@@
    • Notes: Positivity $p(\tau_{i})$ is inversely proportional to negativity $n(\tau_{i})$. So statements like $p(\tau_{i})>3/5$ are equivalent to say the negativity score is $n(\tau_{i})<3/5$.
@@
    • Prompt: “Assess the geographical variability of the texts in the whole web about {topic}. Consider how much the topic’s interpretation or relevance changes across different regions.
    • Abstraction Definition: This property reflects whether the topic is discussed or interpreted differently depending on cultural, legal, or regional factors. High geographical variability means the topic is subject to different norms, laws, or levels of acceptance in various parts of the world.
@@
    • Prompt: “Evaluate the interdisciplinarity of the texts in the whole web about {topic}. Think about how often the topic spans multiple fields or domains (e.g., biology + computer science, philosophy + physics).
    • Abstraction Definition: This measures how frequently the topic draws on or bridges multiple academic or professional fields. An interdisciplinary topic is actively discussed by experts from different areas, each bringing distinct perspectives.
@@
    • Prompt: “Evaluate the time variability of the texts in the whole web about {topic}. Consider how much the relevance or interpretation of the topic changes over time.
    • Abstraction Definition: This refers to how the focus, relevance, or societal perception of the topic shifts over years or decades. A topic with high time variability might be deeply influenced by historical context, cultural trends, or technological advancements.
@@
Notably, we decode the abstraction scores deterministically (T=0T\!=\!0, top-p =0\!=\!0) to minimize sampling variance, so outputs reflect instructions and LLM parameters (i.e., beliefs) rather than randomness.
Appendix B Output Space Abstractions
B.1 Main Features: LLM-Independent Abstraction Functions
The main bias-estimation proxy metrics are:
  1. 1.
    Explanation readability, measured in terms of Gunning Fog index [33].
  2. 2.
    Explanation length, simply measured as character count.
  3. 3.
    Explanation sentiment.
  4. 4.
    Explanation subjectivity.
We estimate the Gunning Fog index using the gunning_fog function from the Python library textstat [54]. Notably, the Gunning Fog index estimates the years of formal education a person needs to understand the text on the first reading. For instance, a fog index of 12 requires the reading level of a United States high school senior (around 18 years old). However, the Fog index is primarily applicable to English and may not accurately reflect readability in other languages. Nonetheless, it can serve as a good, though imperfect, approximation of readability when applied to longer passages or whole documents [55] rather than single sentences, since short texts can distort the score due to sentence length and word choice.
The second metric (explanation length) was implemented using native Python functions for computing characters count in a string.
Finally, sentiment and subjectivity (the last two metrics) were estimated using pre-trained Transformer-based models. The implementation involves text chunking, inference using language models, and aggregation of results.
To analyse text, we use two Hugging Face pipelines:
  • Sentiment Analysis: Uses the “tabularisai/multilingual-sentiment-analysis” model.
  • Subjectivity Classification: Uses the “GroNLP/mdebertav3-subjectivity-multilingual” model.
$\texttt{max\_characters}=\texttt{max\_tokens}\times\texttt{avg\_chars\_per\_token}$ (1)
Each chunk is then processed separately by the respective model.
B.2 Extra Features: LLM-as-a-Judge Protocol
@@
Also in this case, we decode the resulting scores deterministically (T=0T\!=\!0, top-p =0\!=\!0) to minimize sampling variance, so outputs reflect instructions rather than randomness.
@@ 2. Topic 2 label: short explanation of why score {score};
During generation, the sampling temperature was set to T=1T\!=\!1, and top_p was also set to 1 to encourage sufficient diversity in the topics.
Appendix D Bias Injection Procedure
Although the XAI methods we evaluate (e.g., SHAP) rely on input-output analysis, they are model-agnostic and operate without knowledge of the injected rules, they simply perturb inputs and observe output changes. Their low MRR scores (see Appendix H) confirm it is not easy to uncover bias rules.
Bias is injected externally via rule-based system instructions at inference time, not by observing outputs, but by altering the model’s behaviour using handcrafted logic (e.g., based on metadata like commonality or positivity). For instance:
def univar_rule(commonality):
if commonality <= 4/5:
return "Writeoneshortsentence."
return "Writeatleastthreelongparagraphs."
Since ground-truth biases are injected based on LLM abstractions, their existence is guaranteed, making XAI evaluations fully independent of hallucinations and prompt template choices.
Appendix E Examples of LLM-Generated SDG-Related Explanations
@@
Appendix F Examples of XAI-Generated Global Rules
Rule representation.
In our framework, each rule is a conjunctive (when multiple predicates are involved) condition defined over a set of input abstractions, and it characterizes the behaviour of a particular output abstraction. This rule-based representation is a standard format shared by many global rule-extraction algorithms (including the baselines we compare against), and it is intentionally designed to be interpretable by humans.
Rule Impact Direction Importance
$\text{common}\leq 0.5\land\text{positive}>0.5$ Positive 2005.67
$\text{common}>0.30\land\text{common}\leq 0.5\land\text{positive}>0.5\land\text{interdisciplinary}>0.5\land\text{interdisciplinary}\leq 0.70$ Positive 110.85
$\text{common}\leq 0.5\land\text{positive}>0.5\land\text{interdisciplinary}>0.5$ Positive 100.95
$\text{common}>0.5\land\text{positive}\leq 0.70\land\text{interdisciplinary}>0.70\land\text{interdisciplinary}\leq 0.89$ Negative 93.79
$\text{common}>0.30\land\text{common}\leq 0.5\land\text{socially controversial}\leq 0.5\land\text{positive}>0.5\land\text{positive}\leq 0.70$ Positive 59.08
$\text{common}>0.5\land\text{interdisciplinary}\leq 0.5$ Negative 55.70
Table 3: Top-6 rules extracted by RuleSHAP from GPT-4o under $b_{3}$ for the target variable “explanation length.”
Direct interpretability.
These rules can be directly interpreted by humans. For instance, the first rule states:
If a topic is relatively uncommon ($\text{common}\leq 0.5$) and is associated with a high positivity score ($\text{positive}>0.5$), then the model tends to generate longer explanations.
This rule aligns with the injected bias condition $b_{3}$, where explanation length was manipulated based on topic positivity and rarity. The impact direction (positive or negative) indicates whether the rule increases or decreases the value of the target variable (here, explanation length). The importance score reflects the magnitude of this impact and serves to prioritize which rules have the strongest effects.
XAI perspective.
From an XAI standpoint, such rules offer interpretable global explanations of model behaviour:
  • Transparent: each rule is declarative and easy to parse.
  • Diagnostic: they reveal how input abstractions interact to produce biased or unexpected outcomes.
  • Actionable: by showing which abstractions and thresholds drive a behaviour, developers can craft targeted mitigations.
Appendix G SHAP: MRR Scores by Injected Bias Type.
@@ Appendix F Examples of XAI-Generated Global Rules
Table 4: MRR scores for SHAP grouped by injected bias type.
@@
Appendix H RuleSHAP vs. Baselines: Error Analysis
Table 5 reports MRR@1 scores for six different rule-extraction methods (Decision Tree, RuleFit, RuleFit with XGB, RuleSHAP without Step 2, RuleSHAP without Step 3, and RuleSHAP) across five LLMs (GPT-3.5 turbo, GPT-4o mini, GPT-4o, Llama-3.1, and Llama-3.1-70B). The results are grouped by complexity tier and by LLM. Key observations are:
@@
GPT-3.5
turbo
Decision Tree 0 0 0
@@
GPT-4o
mini
Decision Tree 0 0 0
@@
Llama
3.1
Decision Tree 0 0 0
@@
Llama
3.170B
Decision Tree 0 0 0
@@
Appendix I Correlation Certificates